A monopolist’s demand curve is given by P = 100 – 2q.
(a) Find his marginal revenue function.
(b) At what price is marginal revenue zero?
1
Expert's answer
2015-03-10T10:16:27-0400
a. We can construct the marginal revenue function by calculating total
revenue as a function of quantity and then taking the derivative.
TR = P * Q= (100 – 2q)*q = 100q – 2q^2
So, the marginal revenue function can be calculating as follows:
MR = d TR /d Q = 100 – 4q
b. If MR = 0
100 – 4q = 0
4q = 100
q = 25
P = 100 – 2q = 100 – 50 = 50